# Communication-Efficient Soft Actor-Critic Policy Collaboration via Regulated Segment Mixture in Internet of Vehicles

### v2

@@ VI-B Evaluation Metrics
(a) Communication times.
(b) Effective mixture times.
@@ VI-C Simulation Results
TABLE V: Results of Average Reward & Communication Efficiency under “Merge” scenario.
@@
Avg- MAPPO $8$ $4$ $3$ $0.5940\pm 0.0204$ $24,381$ $6.185$ $24,381$ $100\%$
MASAC $0.6584\pm 0.0107$ $43,797$ $11.111$ $43,797$ $100\%$
RSM- MAPPO $0.6173\pm 0.0097$ $24,381$ $6.185$ $10,294$ $42.22\%$
MASAC $1$ $0.6648\pm 0.0065$ $14,599$ $3.704$ $7,938$ $54.37\%$
$5$ $0.6745\pm 0.0120$ $72,995$ $18.518$ $33,484$ $45.87\%$
$7$ $0.6721\pm 0.0072$ $102,102$ $25.902$ $47,122$ $46.15\%$
$3$ $0.6658\pm 0.0063$ $43,797$ $11.111$ $22,661$ $51.74\%$
$1$ $0.6654\pm 0.0048$ $43,797$ $11.111$ $21,024$ $48.00\%$
$2$ $0.6677\pm 0.0064$ $43,797$ $11.111$ $21,083$ $48.14\%$
$6$ $0.6656\pm 0.0059$ $43,797$ $11.111$ $22,128$ $50.52\%$
$72$ $4$ $0.6464\pm 0.0292$ $4,875$ $1.237$ $2,345$ $48.10\%$
$144$ $0.6455\pm 0.0486$ $2,457$ $0.623$ $1,046$ $42.57\%$
@@
Beforehand, we present the performance of Independent SAC (ISAC) and Independent PPO (IPPO) for MARL without any information sharing in Fig. 4, so as to highlight the critical role of inter-agent communication in decentralized cooperative MARL and facilitate subsequent discussions. Fig. 4 reveals that the learning process of these non-cooperative algorithms achieves unstable average reward with greater variance, and suffers from convergence issues even until $300$ epochs. Whereas, when a communication-assisted mixing phase involving the exchange of policy parameters is incorporated, as shown in Fig. 5, the learning process within the multi-agent environment is remarkably enhanced in terms of stability and efficiency. In other words, consistent with our previous argument, integrating DFL into IRL significantly improves training efficiency and ensures learning stability.
@@
In Fig. 5, we also compare the performance of Regulated Segment Mixture (RSM) and that of the direct, unselective Averaging (Avg) method, as mentioned in Section III. It can be observed in Fig. 5, in terms of convergence speed and stability, the simple average mixture method is somewhat inferior. Particularly in more complex scenarios, as in “Merge”, which involves a greater number of RL agents and denser traffic flow compared to the “Figure 8”, the disparity between these two methods becomes more pronounced. The larger variance in the average reward under the average mixture method suggests a more unstable parameter mixing process. Therefore, it validates the effectiveness of RSM and supports the derived theoretical results for selecting useful reference policies and assigning appropriate mixing weights.
On the other hand, Fig. 5 offers a further evaluation of the RSM’s performance on both MASAC and MAPPO. In particular, proposed in our previous work [1], RSM-MAPPO utilizes PPO under traditional MARL framework as the independent local learning algorithm, making it a special case of RSM-MASAC with $\alpha=0$ and a different policy advantage estimation under our more general reanalyses. Evidently, in both scenarios, the SAC-based algorithm demonstrates faster convergence and overall superior performance compared to the PPO-based algorithm under the same learning rate. We believe this is primarily attributed to the differences in exploration mechanism and sample efficiency. Specifically, benefiting from the introduced entropy item under MERL framework, SAC encourages policies to better explore environments, thus capably avoiding local optima and potentially learning faster.
(a) the number of segments $P$
(b) the number of replicas $\kappa$
@@
Figure 8: The performance of RSM-MASAC with respect to different hyperparameters ((a) the number of segments $P$ and (b) the number of replicas $\kappa$), based on the testing episodes after $150$, $160$, $\cdots$, $250$ training epochs under “Merge”.
@@
Next, we evaluate the impact of regulated segmentation. Fig. 7 shows that the incorporation of segmentation (i.e., $P=4$) also leads to an improvement in the successful mixing rate $\rho_{\text{r}}$ of the policy, apart from better utilizing the available bandwidth as discussed in Section V-B. The improvement lies in the introduction of a certain level of randomness, consistent with the idea to encourage exploration by adding an entropy term in MERL. But this randomness is also regulated by the policy parameter mixing theorem in Theorem 2 without compromising performance. To embody the selection process of reconstructed referential policies, we further present the heatmap comparison between communication times and successful mixture times in Fig. 7. It can be observed from Fig. 7(a) that vehicles communicate frequently with their neighboring vehicles, but since not all communication packets (i.e., model segments) are utilized for improving local policy performance, it leads to the uneven and asymmetrical heatmap for the successful mixture of segments in Fig. 7(b). Our derived performance improvement bound and practical parameter mixture metric regulate these segments, demonstrating that the data from different agents contribute variably to the mixture.
In Fig. 8, we further discuss the effects of the number of segments $P$ and replicas $\kappa$ under RSM-MASAC. Considering that RSM-MASAC generally converges from $250$ to $300$ testing epochs, the results in the boxplots of Fig. 8 are derived from testing epochs ranging from $150$ to $250$, during which RSM-MASAC experiences more fluctuation and slower convergence, so as to better illustrate the performance during training. Specifically, Fig. 8(a) displays the distribution of average reward, where each box represents the InterQuartile Range (IQR). The lower and upper edges of each box denote the first and third quartiles, respectively. Besides, the line inside the box signifies the median average reward for each segment number. Observing Fig. 8(a), a trend emerges showing that the median average reward subtly increases with the growing number of segments. Besides, Fig. 8(b) shows the median average reward slightly increases along with larger $\kappa$ values, suggesting a trend that more replicas may produce with a higher reward or a higher probability of exploring a better policy, at some cost of a heavier perturbation of the learning process.
@@
In addition, we take the average reward within $250-300$ testing epochs as the final converged performance, and summarize corresponding experimental results and details about communication overhead in Table V, which takes the V2V communication examples’ settings in V-B. We can find that with $P$ increasing, the mixing rate $\rho_{r}$ also increases first, but slightly decreases when $P=6$, since the aggregation target of reconstructed policy parameters for an over-large $P$ might be mottled and lose the integrity. Besides, a larger $\kappa$ gives the increase of the communication overhead $\psi$, while the final performance gains, especially for $\kappa=5$, $\kappa=7$, are limited, due to that the algorithm has converged to a relatively good policy as in centralized FMARL [27].
@@
In addition, we conduct more ablation studies to evaluate the design of $Q$ network on the learning process. In Fig. 10, we analyze the impact of different target smoothing coefficients $\varrho$ on the learning process. It can be observed that a larger value of the target smoothing coefficient, such as $10^{-2}$, can significantly accelerate the learning, but it also introduces disturbances, leading to non-steady learning. On the other hand, a smaller value like $10^{-4}$ or $10^{-5}$ noticeably slows down the learning speed, affecting the performance of the algorithm. Hence, we take $\varrho=10^{-3}$ as the default value. Moreover, in Fig. 10, we also evaluate the adoption of dual $Q$ networks. It can be observed that employing dual $Q$ networks indeed accelerates the learning process.
Furthermore, we also discuss the impact of communication intervals on the learning process, as shown in Fig. 11. It is evident that larger communication intervals result in reduced learning speeds and significant instability in the learning process.
@@
Meanwhile, we also investigate the effect of model fine-tuning in decentralized IoV speed control settings. Specifically, in an environment full of $14$ DRL-controlled CAVs, we compare the performance between a continually fine-tuning policy and a well-trained, convergent policy (i.e., RSM-MASAC in Fig. 5) to execute inference only (i.e., test only in DRL description). The results in Fig. 13 show that the fine-tuned policy in new conditions can reduce the occurrence of collisions at the beginning and continuously achieve better performance than just testing. Furthermore, to investigate the impact of possible packet loss when communicating on V2V links, we provide the result for different Packet Reception Rate (PRR) in Fig. 13. It can be seen that a packet loss of less than $20\%$ has a trivial effect on performance, demonstrating that DNN parameter transmission in VFL can yield good robustness in most real communication environments.
VII Conclusions
@@
In this paper, we have proposed a communication-efficient algorithm RSM-MASAC as a promising solution to enhance communication efficiency and policy collaboration in distributed MARL, particularly in the context of dynamic IoV environments. By delving into the policy parameter mixture function, RSM-MASAC has provided a novel means to leverage and boost the effectiveness of distributed multi-agent collaboration. In particular, RSM-MASAC has successfully transformed the classical means of complete parameter exchange into segment-based request and response, which significantly facilitates the construction of multiple referential polices and simultaneously captures enhanced learning diversity. Moreover, in order to avoid performance-harmful parameter mixture, RSM-MASAC has leveraged a theory-established regulated mixture metric, and selects the contributive referential policies with positive relative policy advantage only. Finally, extensive simulations in the mixed-autonomy traffic control scenarios have demonstrated the effectiveness of the proposed approach.
@@ VII-A Proof of Soft Policy Improvement
$=\arg\min_{\pi\in\prod}J_{\pi_{\text{old}}}(\pi(\cdot|s))$
@@
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s},{a})-\log\pi_{\text{old}}(a|s)-\log Z^{\pi_{\text{old}}}(s)\right]$
@@
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}({s},{a})\right]+\alpha H(\pi_{\text{old}}(\cdot|{s}))=V_{\pi_{\text{old}}}(s)$
@@ VII-B Proof of Theorem 1
$\mathbb{E}_{a\sim\pi}\left[A_{\pi}(s,a)\right]=-\alpha H(\pi(\cdot|s))$
@@
$\mathop{\mathbb{E}}\limits_{a\sim\pi_{\text{mix}}}\left[A_{\pi}(s,a)\right]=\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s))$
@@
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}$ $\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s))$
@@
$=$ $\eta(\pi_{\text{mix}})-\eta(\pi)-\mathbb{E}_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\alpha H(\pi_{\text{mix}}(\cdot|s_{t}))]\right]$
@@
$\qquad+(1-\beta)H(\pi(\cdot|s))$
@@
$=$ $\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))+\beta H(\tilde{\pi}(\cdot|s))+(1-\beta)H(\pi(\cdot|s))$
@@
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[D_{\text{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
@@
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$
@@
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\leq\varepsilon$
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}\geq 1;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\geq-\varepsilon$
@@
$=$ $\beta\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi)}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\!-\!2\beta\rho_{t}\varepsilon$

### v3

@@ VI-B Evaluation Metrics
(a) Communication times
(b) Effective mixture times
@@ VI-C Simulation Results
TABLE V: Results of Average Reward & Communication Efficiency under “Merge”.
@@
Avg- MAPPO $8$ $4$ $3$ $0.5351\pm 0.0031$ $1,626$ $0.412$ $1,626$ $100\%$
MASAC $0.4818\pm 0.0161$ $29,270$ $7.425$ $29,270$ $100\%$
RSM- MAPPO $0.6528\pm 0.0199$ $1,546$ $0.392$ $711$ $45.99\%$
MASAC $1$ $0.6538\pm 0.0250$ $10,223$ $2.593$ $5,213$ $50.99\%$
$5$ $0.6636\pm 0.0115$ $41,333$ $10.486$ $20,410$ $49.38\%$
$7$ $0.6574\pm 0.0156$ $38,120$ $9.670$ $19,722$ $51.74\%$
$3$ $0.6614\pm 0.0165$ $30,486$ $7.734$ $13,827$ $45.36\%$
$1$ $0.6535\pm 0.0253$ $30,066$ $7.627$ $12,117$ $40.30\%$
$2$ $0.6206\pm 0.0793$ $29,634$ $7.518$ $14,549$ $49.10\%$
$6$ $0.6505\pm 0.0282$ $30,126$ $7.643$ $13,902$ $46.15\%$
$72$ $4$ $0.6712\pm 0.0282$ $2,351$ $0.596$ $1,357$ $57.72\%$
$144$ $0.6576\pm 0.0269$ $1,227$ $0.311$ $708$ $57.70\%$
@@
Beforehand, we present the performance of Independent SAC (ISAC) and Independent PPO (IPPO) for MARL without any information sharing in Fig. 5, so as to highlight the critical role of inter-agent communication in decentralized cooperative MARL and facilitate subsequent discussions. Fig. 5 reveals that the learning process of these non-cooperative algorithms achieves unstable average reward with greater variance, and suffers from convergence issues. Whereas, when a communication-assisted mixing phase involving the exchange of policy parameters is incorporated, as shown in Fig. 6, the learning process within the multi-agent environment is remarkably enhanced in terms of stability and efficiency. In other words, consistent with our previous argument, integrating DFL into IRL significantly improves training efficiency and ensures learning stability.
@@
In Fig. 6, we also compare the performance of Regulated Segment Mixture (RSM) and that of the direct, unselective Averaging (Avg) method, as mentioned in Section III. Note that to speed up the calculation, we use a batch average gradient to approximate the calculation of $F(\theta)$ here. In Fig. 6, it can be observed that in terms of convergence speed and stability, the simple average mixture method is somewhat inferior. Particularly in more complex scenarios, as in “Merge”, which involves a greater number of RL agents and denser traffic flow compared to the “Figure 8”, the disparity between these two methods becomes more pronounced. The larger variance in the average reward under the average mixture method suggests a more unstable parameter mixing process, while the improvement in reward value implies the efficiency of communication information sharing. Therefore, it validates the effectiveness of RSM and supports the derived theoretical results for selecting useful reference policies and assigning appropriate mixing weights.
(a) predefined number of segments $P$
(b) predefined number of replicas $\kappa$
@@
Figure 9: The performance of RSM-MASAC with respect to different hyperparameters ((a) predefined number of segments $P$ and (b) predefined number of replicas $\kappa$), based on the testing episodes after $130$, $140$, $\cdots$, $200$ training epochs under “Merge”.
@@
On the other hand, Fig. 6 offers a further evaluation of the RSM’s performance on both MASAC and MAPPO under the same policy DNN structure, learning rate and gradient clipping condition. In particular, as proposed in our previous work [1], RSM-MAPPO utilizes PPO, which features lower sampling efficiency and policy update frequency, as the independent local learning algorithm under the traditional MARL framework, making it a special case of RSM-MASAC with $\alpha=0$ and a different policy advantage estimation under our more general reanalyses. Evidently, in both scenarios, RSM-MASAC demonstrates faster convergence and overall superior performance compared to RSM-MAPPO. We believe this is primarily attributed to the differences in exploration mechanism and sample efficiency. Specifically, benefiting from the introduced entropy item under MERL framework, SAC encourages policies to better explore environments, thus capably avoiding local optima and potentially learning faster.
Next, we evaluate the impact of regulated segmentation. Fig. 8 shows that the incorporation of segmentation (i.e., $P=4$) also leads to an improvement in the successful mixing rate $\rho_{\text{r}}$ of the policy, apart from better utilizing the available bandwidth as discussed in Section V-B. The improvement lies in the introduction of a certain level of randomness, consistent with the idea to encourage exploration by adding an entropy term in MERL. But this randomness is also regulated by the policy parameter mixing theorem in Theorem 2 without compromising performance. To embody the selection process of reconstructed referential policies, we further present the heatmap comparison between communication times and successful mixture times in Fig. 8. It can be observed from Fig. 8(a) that agents communicate frequently with their neighboring agents, but since not all communication packets (i.e., model segments) are utilized for improving local policy performance, it leads to the uneven and asymmetrical heatmap for the successful mixture of segments in Fig. 8(b). Our derived performance improvement bound and practical parameter mixture metric regulate these segments, demonstrating that the data from different agents contribute variably to the mixture.
In Fig. 9, we further examine the impact of varying the predefined segmentation granularity $P$ and replicas $\kappa$ under RSM-MASAC. Notably, for each agent, the actual segment number and replicas at each time are also affected by the number of neighbors in the communication range, as addressed in Section V-A. The boxplots display the distribution of average reward collected during testing epochs ranging from $130$ to $200$, a period marked by increased fluctuations and slower convergence. These outcomes are inherently subject to the stochastic nature of the environment interacting with the RL agents and here we only indicate a rough trend that the median average reward shows a modest increase with a greater number of segments in Fig. 9(a), and a slight upward trend with an increase replicas in Fig. 9(b). However, these variations are minimal and contingent upon specific environmental conditions that are not consistent, thus warrant further discussions to balance communication bandwidth and computational overhead in practical applications.
@@
In addition, we take the average reward within the last five testing epochs as the final converged performance, and summarize corresponding results and details about communication overhead in Table V, which takes the V2V communication examples’ settings in V-B.
@@
In addition, we conduct more ablation studies to evaluate the design of $Q$ network on the learning process. In Fig. 11, we analyze the impact of different target smoothing coefficients $\varrho$ on the double $Q$-learning process. It can be observed that a larger value of the target smoothing coefficient $\varrho$, such as $10^{-2}$, can significantly accelerate the learning, but it also introduces disturbances, leading to non-steady learning. On the other hand, a smaller value like $10^{-4}$ or $10^{-5}$ noticeably slows down the learning speed and also introduces instability, affecting the performance of the algorithm. Moreover, in Fig. 11, we also evaluate the adoption of dual $Q$ networks. It can be observed that employing dual $Q$ networks indeed accelerates the learning process. Furthermore, we also discuss the impact of communication intervals on the learning process, as shown in Fig. 12. It is evident that larger communication intervals result in reduced learning speeds and significant instability in the learning process, the same as classical FL processes.
@@
Meanwhile, we also investigate the effect of model fine-tuning in decentralized IoV speed control settings. Specifically, in a new environment full of $14$ DRL-controlled CAVs, we compare the performance between a continually fine-tuning policy and a well-trained, convergent policy (i.e., RSM-MASAC in Fig. 6) to execute inference only (i.e., test only in DRL description). The results in Fig. 14 show that the fine-tuned policy in new conditions can reduce the occurrence of collisions at the beginning and continuously achieve better performance than just testing. Furthermore, to investigate the impact of possible packet loss when communicating on D2D or V2V links, we provide the result for different Packet Reception Rate (PRR) in Fig. 14. It can be seen that a packet loss of less than $40\%$ primarily affects the convergence speed at the beginning, but produces trivial differences on the converged performance, demonstrating that DNN parameter transmission can yield good robustness in most real communication environments.
VII Conclusions & Future Work
@@
In this paper, we have proposed a communication-efficient algorithm RSM-MASAC as a promising solution to enhance communication efficiency and policy collaboration in distributed MARL, particularly in the context of highly dynamic environments. By delving into the policy parameter mixture function, RSM-MASAC has provided a novel means to leverage and boost the effectiveness of distributed multi-agent collaboration. In particular, RSM-MASAC has successfully transformed the classical means of complete parameter exchange into segment-based request and response, which significantly facilitates the construction of multiple referential policies and simultaneously captures enhanced learning diversity. Moreover, in order to avoid performance-harmful parameter mixture, RSM-MASAC has leveraged a theory-established regulated mixture metric, and selects the contributive referential policies with positive relative policy advantage only. Finally, extensive simulations in the mixed-autonomy traffic control scenarios have demonstrated the effectiveness of the proposed approach. Notably, we use an approximate calculation here to reduce the computation complexity and speed up the calculation of $F(\theta)$, thus supporting the validation of the mixed performance improvement bound theorem. Despite its efficiency, its impact is still under investigation, and other methods to simplify FIM computations can be explored and substituted in the future.
@@ VII-A Proof of Soft Policy Improvement
$=\arg\min_{\pi\in\prod}J_{\pi_{\text{old}}}(\pi(\cdot|s)){.}$
@@
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[\frac{1}{\alpha}Q^{\pi_{\text{old}}}({s},{a})-\log\pi_{\text{old}}(a|s)-\log Z^{\pi_{\text{old}}}(s)\right]{.}$
@@
$\geq\mathbb{E}_{{a}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}({s},{a})\right]+\alpha H(\pi_{\text{old}}(\cdot|{s}))=V_{\pi_{\text{old}}}(s){.}$
@@ VII-B Proof of Theorem 1
$\mathbb{E}_{a\sim\pi}\left[A_{\pi}(s,a)\right]=-\alpha H(\pi(\cdot|s)){.}$
@@
$\mathop{\mathbb{E}}\limits_{a\sim\pi_{\text{mix}}}\left[A_{\pi}(s,a)\right]=\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s)){.}$
@@
=$\stackrel{{\scriptstyle\mathrm{(b)}}}{{=}}$ $\beta\mathop{\mathbb{E}}\limits_{a\sim\tilde{\pi}}\left[A_{\pi}(s,a)\right]-(1-\beta)\alpha H(\pi(\cdot|s)){,}$
@@
$=$ $\sum_{t=0}^{\infty}\gamma^{t}\left[\mathbb{E}_{\begin{subarray}{c}s\sim P(s_{t};\pi_{\text{mix}})\\ a\sim\pi_{\text{mix}}\end{subarray}}[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))]\right]{.}$
@@
$=$ $\eta(\pi_{\text{mix}})-\eta(\pi)-\mathbb{E}_{\tau_{0}\sim\pi_{\text{mix}}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\alpha H(\pi_{\text{mix}}(\cdot|s_{t}))]\right]{.}$
@@
$=$ $\sum_{t=0}^{\infty}\gamma^{t}\left[\mathbb{E}_{\begin{subarray}{c}s\sim P(s_{t};\pi_{\text{mix}})\\ a\sim\pi_{\text{mix}}\end{subarray}}[A_{\pi}(s,a)+\alpha H(\pi_{\text{mix}}(\cdot|s))]\right]{.}$
@@
$\qquad+(1-\beta)H(\pi(\cdot|s)){.}$
@@
$=$ $\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))+\beta H(\tilde{\pi}(\cdot|s))+(1-\beta)H(\pi(\cdot|s)){.}$
@@
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[D_{\text{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]{,}$
@@
$\quad+\alpha\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi_{\text{mix}})}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]{,}$
@@
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}=0;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\leq\varepsilon{;}$
$\mathop{\mathbb{E}}\limits_{s\sim P(s_{t}|c_{t}\geq 1;\pi_{\text{mix}})}\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\geq-\varepsilon{.}$
@@
$=$ $\beta\!\mathop{\mathbb{E}}\limits_{s\sim P(s_{t};\pi)}\!\left[\sum_{a}\tilde{\pi}(a|s)\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]\right]\!-\!2\beta\rho_{t}\varepsilon{,}$
@@
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]{.}$